OpenStudy (aaronandyson):

Convert 110 degrees and 30 minutes to radians.

OpenStudy (smoothini):

1.92858882 radians

OpenStudy (aaronandyson):

Please show the working.

OpenStudy (xapproachesinfinity):

110 degree into radians 110 times pi/180

OpenStudy (xapproachesinfinity):

now for minutes 30 minutes 2pi/60 minutes then add up first and second results to get the answer

OpenStudy (aaronandyson):

110.5 times pi/180?

OpenStudy (xapproachesinfinity):

are you asking why i should multiply 110 by pi/180

OpenStudy (aaronandyson):

I think that 110.5 time pi/180 is the final answer.

OpenStudy (xapproachesinfinity):

well i didn't do the calculations

OpenStudy (xapproachesinfinity):

\[\frac{11\pi}{18}+30\times \frac{2\pi}{60}\] \[\frac{11\pi}{18}+\pi=\frac{29\pi}{18}\]

OpenStudy (xapproachesinfinity):

that's the answer

OpenStudy (xapproachesinfinity):

you get it?